OpenStudy (anonymous):

Can someone explain how to do this? |x - 6| > 2

OpenStudy (anonymous):

i can

OpenStudy (anonymous):

(I think you're supposed to do it now, jasonsmart ;))

OpenStudy (anonymous):

Then please do explain.

OpenStudy (anonymous):

oh I GOT THIS!!

OpenStudy (slaaibak):

|x| > a means x > a or x<-a where a is nonnegative

OpenStudy (anonymous):

so what you do .......YOU SON A MOTHER!!!!

OpenStudy (anonymous):

I did not understand a thing....

OpenStudy (anonymous):

yes my time to shine!

OpenStudy (anonymous):

al right so what you do is add six to the other side of the greater sign and it will be X>2+6 so the answer would be X<8

OpenStudy (anonymous):

4<x V x>8

OpenStudy (anonymous):

....

OpenStudy (anonymous):

jason I think you're wrong. If the solution is x<8 then, for example |6-6|>2 is not true. My solution is right!

OpenStudy (anonymous):

whut....

OpenStudy (anonymous):

Okay... Come on, I need to know how to solve these equations... So, I need a correct explanation.

OpenStudy (anonymous):

.....you dare condescend me!.....

OpenStudy (anonymous):

.... well you didnt give a very god question

OpenStudy (anonymous):

|x-6|>2 You have to do x-6>2 and x-6<-2 First is x>8 and second is x<4 Right?

OpenStudy (anonymous):

Achab is right.

OpenStudy (anonymous):

whut....?

OpenStudy (anonymous):

I don't understand his explanation....

OpenStudy (anonymous):

I'll explain you now.

OpenStudy (anonymous):

....

OpenStudy (anonymous):

\[\Large x \in ( - \infty ,4) \cup (8,\infty )\]

OpenStudy (anonymous):

Basically TheHero, the absolute value of a number is the positive form of that number. |-4| = 4, |-199123| = 199123, |12| = 12 and so on. This question means that if you take 6 away from x, then the absolute result must be greater than two. If x is greater than 4 but less than 8 - say, 5 - then you'll end up with |5-6| = 1, which is not greater than 2, so it is invalid. But if you take a value less than 4 - say, 2 - then you end up with |2-6| = 4 which is greater than 2 so is valid. Similarly, over 8 - say, 15 - you get |15-6| = 9 which is greater than 2 and also valid.

OpenStudy (anonymous):

we are solving for x?

OpenStudy (anonymous):

I think i'm better off posting the questions.... Thanks anyways...

OpenStudy (anonymous):

sorry i tried...

OpenStudy (anonymous):

It's okay~ , you did good c:

OpenStudy (anonymous):

yaey

OpenStudy (anonymous):

\[|x-6|>2\] means that you have to search which x, subtracted from 6 and Absolutized, are higher than 2 (2,01; 2,00000001; 3;4, etc.). You have to do 2 disequations. The first one you have to do: x-6>2. Only remove the | | and solve it. It will be x>8, right? Well, for the second you have to do: x-6<-2 You have to change the > to < and change the 2 to -2 (change the sign). Now x<4 right? Well, in the first you found x>8, in the second x<4, then, the xs are under 4 and higher than 8, right? The solution is: \[x<4 \] V (and) \[x>8\]. Prove yourself that if you take a x lower than 4 or higher than 8 it'll be correct. Hope you've understand!